OpenStudy (anonymous):

[Help Me for Medal And Fan :3] If the perimeter of one face of a cube is 8cm, then its total surface area is: a) 16cm^2 b) 24cm^2 c) 48cm^2

OpenStudy (anonymous):

@ganeshie8 Plz Help me, bro?

ganeshie8 (ganeshie8):

face of a cube is a square so the side of cube would be 8/4 = 2

ganeshie8 (ganeshie8):

how many faces are there in a cube ?

OpenStudy (anonymous):

6 @ganeshie8

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The surface area of a cube is 6a^2

OpenStudy (anonymous):

You already have a (look at ganeshies work)

ganeshie8 (ganeshie8):

find the area of one face : \[2\times 2 = 4\] multiply that by \(6\) to get the area of all faces \[4\times 6 = 24\]

OpenStudy (anonymous):

 this is the perimeter, it's the distance around the square, just giving a visual if you don't know it.

OpenStudy (anonymous):

Thanks guys, Wish I could medal you both, I have fanned both tho. :) @iambatman @ganeshie8

OpenStudy (anonymous):

Np :)